The Grade 11 students of Capiz National High School took a 100 item Statistics exam. The scores fit the normal distribution with a mean of 75 and a standard deviation of 10. What is the percentage of students that got a score of 80 to 85?

1
Expert's answer
2022-04-17T17:12:50-0400

We have a normal distribution, "\mu=75, \sigma=10."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"z_1=\cfrac{80-75}{10}=0.5;\\
z_2=\cfrac{85-75}{10}=1;\\
P(80<X<85)=P(0.5<Z<1)=\\
=P(Z<1)-P(Z<0.5)=\\
=0.8413-0.6915=\\
=0.1498=14.98\% \\
\text{ (from z-table).}"